### v3

# MML Probabilistic Principal Component Analysis

Abstract
Principal component analysis (PCA) is perhaps the most widely used method for data dimensionality reduction. A key question in PCA is deciding how many factors to retain. This manuscript describes a new approach to automatically selecting the number of principal components based on the Bayesian minimum message length method of inductive inference. We derive a new estimate of the isotropic residual variance and demonstrate that it improves on the usual maximum likelihood approach. We also discuss extending this approach to finite mixture models of principal component analyzers.
Index Terms:
Principal component analysis, minimum message length, bias, model selection.
I Introduction
The principal component analysis (PCA) model [1] postulates that $N$ independent realisations of $K$-dimensional data ${\bf x}_{i}\in\mathbb{R}^{K}$ ($i=1,\ldots,N$) are described as
${\bf x}_{i}=v_{i1}{\bf a}_{1}+\cdots+v_{iJ}{\bf a}_{J}+\bm{\epsilon}_{i}=\left(\sum_{j=1}^{J}v_{ij}{\bf a}_{j}\right)+\bm{\epsilon}_{i},\quad\bm{\epsilon}_{i}\sim{\rm N}({\bf 0}_{K},\sigma^{2}{\bf I}_{K}),$ (1)
where $\{{\bf a}_{1},\ldots,{\bf a}_{J}\}$ are the $J(<K)$ latent (unobserved) factor loadings with each factor loading ${\bf a}_{j}\in\mathbb{R}^{K}$, and ${\bf v}_{ij}\sim{\rm N}(0,1)$ are the factor scores distributed as per the standard normal distribution. It is assumed that the residuals follow an isotropic zero mean normal distribution with the variance-covariance matrix $\sigma^{2}{\bf I}_{K}$. We can write this PCA model in matrix notation
${\bf x}_{i}={\bf A}{\bf v}_{i}+\bm{\epsilon}_{i},\quad{\bf A}\in\mathbb{R}^{K\times J},\quad{\bf v}_{i}\in\mathbb{R}^{J},\quad\bm{\epsilon}_{i}\sim N({\bf 0},\sigma^{2}{\bf I}_{K}),$ (2)
where $i=(1,\ldots,N)$, ${\bf A}=({\bf a}_{1},\ldots,{\bf a}_{J})$ and ${\bf V}=({\bf v}_{1},\ldots,{\bf v}_{N})\in\mathbb{R}^{J\times N}$. Integrating out the factor scores yields the multivariate Gaussian marginal distribution of the data
${\bf x}_{i}\sim N({\bf 0}_{K},\bm{\Sigma}),\quad\bm{\Sigma}={\bf A}{\bf A}^{\prime}+\sigma^{2}{\bf I}_{K}.$ (3)
This setup is also known as the classical spiked covariance model where the covariance matrix $\bm{\Sigma}$ has $J$ large population eigenvalues $\lambda_{1}>\lambda_{2}>\cdots>\lambda_{J}$ (the spikes) that represent strong data signals with $\lambda_{j}=\alpha_{j}^{2}+\sigma^{2}$, while the remaining (KJ)(K-J) population eigenvalues $\lambda_{J+1}=\lambda_{J+2}=\cdots=\lambda_{K}=\sigma^{2}$ are small and represent noise.
The probabilistic principal component model suffers from identifiability constraints [2, 3]. A key reason for this is that the latent factors affect the likelihood function only through their outer product ${\bf A}{\bf A}^{\prime}$, which implies that an estimate of the factors can only be determined up to a rotation. To ensure that the matrix ${\bf A}{\bf A}^{\prime}$ is identifiable asymptotically, the eigenvalues of ${\bf A}^{\prime}{\bf A}$ must be uniformly bounded away from both zero and infinity as $N\to\infty$; this is known as the pervasive assumption (see, for example, [4]). Additionally, to ensure the PCA model is not overparameterised, the maximum number of latent factors to be estimated cannot exceed
$J_{\text{MAX}}\leq K+\frac{1}{2}\left(1-\sqrt{8K+1}\right),$ (4)
see [5] (pp. 108) for details. For example, when $K=4,5,6$ we have $J_{\text{MAX}}=1,2,3$, respectively. Tipping and Bishop [6] showed how to interpret standard PCA model in a probabilistic framework and obtained maximum likelihood estimates of the latent factors and residual variance.
There exists a large volume of literature on PCA (e.g., [7]), and Bayesian PCA (e.g., [8, 9, 10]) models. An important decision for effective PCA is estimating how many principal components should be included in the model (see, for example, [11, 12, 13, 14, 15]). Retaining only a few principal components may result in a loss of information while using more principal components than necessary will weaken the overall signal strength. If the sample size is large, or we consider the asymptotic regime as $N\to\infty$ with $K$ fixed, the eigenvalues of the sample covariance matrix $\delta_{j}$ converge almost surely to the population eigenvalues, $\delta_{j}\xrightarrow{\text{a.s.}}\lambda_{j}$. The noise eigenvalues of the sample covariance matrix converge to the same residual variance $\sigma^{2}$ with probability one. In contrast, the $j$-th signal eigenvalue converges to $(\sigma^{2}+\lambda_{j})$ with probability one. However, when the sample size is small to moderate, the sample noise eigenvalues tend to have large variance and can be significantly different from each other (see, for example, [16]).
The current approaches to estimating the number of principal components can broadly be divided into three categories [17]: (i) model selection criteria, (ii) the scree plot, and (iii) thresholding based on random matrix theory. To select the number of principal components, model selection criteria generally minimise the negative log-likelihood function subject to a penalty on the model complexity. The approach introduced in this manuscript fits into this category. Other examples include the commonly used Akaike’s information criterion (AIC) and Bayesian information criterion (BIC) [11, 13], as well as improved variants thereof such as the generalised information criterion [14] and normalized maximum likelihood [18, 19]. Methods based on the scree plot estimate the number of principal components by visual inspection (i.e., by looking for an ‘elbow’ in the plot of sorted eigenvalues of the sample correlation matrix) or the corresponding test statistics [20]. Lastly, methods based on random matrix theory estimate the number of principal components by thresholding the eigenvalues of the sample covariance matrix, where the threshold is selected based on random matrix theory results [21, 22, 15].
This manuscript examines the estimation of the probabilistic PCA model under the Bayesian minimum message length (MML) inductive inference framework. We develop a new model selection criterion that automatically determines the number of principal components that should be retained as well as a new estimate for the residual variance that improves upon the standard maximum likelihood estimate. Although single and multiple factor analysis has been examined within the MML framework by [23] and [24] respectively, this manuscript departs from the earlier work in the following:
  • We consider the marginal distribution of the data (3) rather than the model (1) analysed by [24].
  • Using polar decomposition, we write the factor load matrix $\bf A$} as a product of an orthogonal matrix and a diagonal matrix representing the direction and length of the loadings, respectively. Unlike earlier MML approaches, we parameterize the orthogonal matrix via Givens rotations to explicitly capture orthogonality constraints.
  • We use matrix polar decomposition to develop a prior distribution for the latent factors $\bf A$} that is a product of a matrix variate Cauchy distribution and a uniform distribution over the corresponding Stiefel manifold.
  • We obtain analytic MML estimates of the parameters and find a polynomial whose roots yield the MML estimate of the residual variance.
  • We characterise the bias of the MML estimate of residual variance and show that it improve on the corresponding maximum likelihood estimate by a factor approximately proportional to $K$.
  • We show that the MML threshold for detecting a latent factor agrees with the Baik-Ben Arous-Péché (BBP) phase transition threshold [25]. The MML threshold is slightly higher than the theoretical distinguishability limit to prevent false positives caused by finite-sample fluctuations of the residual variance estimate.
II Maximum likelihood estimation
This section summarises the results of [6]. The negative log-likelihood of the data under the probabilistic PCA model (3) is
$\ell(\bm{\theta})=\frac{NK}{2}\log(2\pi)+\frac{N}{2}\log|\bm{\Sigma}|+\frac{N}{2}{\rm tr}\left(\bm{\Sigma}^{-1}{\bf S}_{x}\right)$ (5)
where ${\bf S}_{x}=\frac{1}{N}\sum_{i}{\bf x}_{i}{\bf x}_{i}^{\prime}$ is the sample variance-covariance matrix. We have the observed data $\bf X$} and wish to estimate the number of latent factors $J$ and all parameters $\bm{\theta}=\left\{{\bf A},\sigma^{2}\right\}$. Differentiating the negative log-likelihood with respect to the factor loads
$\partial\ell(\bm{\theta})$ $=N{\rm tr}\,{\bf A}^{\prime}\bm{\Sigma}^{-1}(\partial{\bf A})-N{\rm tr}\left({\bf A}^{\prime}\bm{\Sigma}^{-1}{\bf S}_{x}\bm{\Sigma}^{-1}(\partial{\bf A})\right)$
and setting the derivatives to zero we get
${\bf S}_{x}{\bf\Sigma}^{-1}{\bf A}$ $={\bf A}$
Consider the singular value decomposition ${\bf A}={\bf U}{\bf L}{\bf V}^{\prime}$, where ${\bf U}\in\mathbb{R}^{K\times J}$, ${\bf L}={\rm diag}(\lambda_{1},\ldots,\lambda_{j})$ and ${\bf V}\in\mathbb{R}^{J\times J}$ is an orthogonal matrix. Noting that $\bm{\Sigma}^{-1}{\bf A}={\bf U}{\bf L}({\bf L}^{2}+\sigma^{2}{\bf I}_{J})^{-1}{\bf V}^{\prime}$, we have
${\bf S}_{x}{\bf U}$ $={\bf U}({\bf L}^{2}+\sigma^{2}{\bf I}_{J})$
${\bf S}_{x}{\bf u}_{j}$ $=(\lambda_{j}^{2}+\sigma^{2}){\bf u}_{j},\quad(j=1,\ldots,J),$
which is an example of the eigenvalue problem. That is, $\bf U$} is a $(K\times J)$ matrix whose columns are the top $J$ eigenvectors of the sample covariance matrix ${\bf S}_{x}$ corresponding to the $J$ largest eigenvalues
$\delta_{j}=\lambda^{2}_{j}+\sigma^{2},\quad j=1,\ldots,J,$ (6)
where $\lambda_{j}=(\delta_{j}-\sigma^{2})^{\frac{1}{2}}$ is the $j$-th largest singular value of $\bf A$}. Without loss of generality we assume that $\delta_{1}>\delta_{2}>\ldots>\delta_{K}>0$ throughout the manuscript. This implies that the maximum likelihood estimate is
$\hat{\bf A}_{\text{ML}}={\bf U}(\bm{\Delta}-\sigma^{2}{\bf I}_{J})^{\frac{1}{2}}{\bf O},\quad\bm{\Delta}={\rm diag}(\delta_{1},\ldots,\delta_{J})$ (7)
where $\bf O$} is an arbitrary (orthogonal) rotation matrix and $\bm{\Delta}$ is a diagonal matrix with the $J$-th largest eigenvalues of ${\bf S}_{x}$. Substituting the maximum likelihood estimate of the factor loads into the negative log-likelihood we have
$\ell(\sigma,\hat{\bf A}_{\text{ML}})=\frac{NK}{2}\log(2\pi)+\frac{N}{2}\sum_{i=1}^{J}\log\delta_{j}+\frac{N(K-J)}{2}\log\sigma^{2}+\frac{NJ}{2}+\frac{N}{2\sigma^{2}}\sum_{j=J+1}^{K}\delta_{j}.$ (8)
The concentrated negative log-likelihood is minimised by
$\hat{\sigma}^{2}_{\rm ML}=\frac{1}{K-J}\sum_{j=J+1}^{K}\delta_{j}$ (9)
which is the empirical average of the (KJ)(K-J) smallest eigenvalues of the sample variance-covariance matrix. Tipping and Bishop [6] show that these estimates minimise the negative log-likelihood and discuss other saddle points of the log-likelihood function.
III Minimum message length analysis of the PCA model
The minimum message length (MML) principle [26, 27, 28, 29] of inductive inference is based on ideas from information theory, Bayesian statistics and data compression. MML considers the standard tasks of parameter estimation and model selection as data compression problems. Given data ${\bf x}\in\mathcal{X}$, the key idea behind MML is to compute the minimum length of a message that describes the data. The MML message by design encodes both a model for the data as well as the data itself, and must be decodable by a receiver who does not know the data. The two parts of an MML message are:
  1. 1.
    the assertion: describes the structure of the model, including all model parameters $\bm{\theta}\in\bm{\Theta}\in\mathbb{R}^{P}$. Let $I(\bm{\theta})$ denote the codelength of the assertion.
  2. 2.
    the detail: describes the data ${\bf x}\in\mathcal{X}$ using the model $p({\bf x}|\bm{\theta})$ nominated in the assertion. Let $I({\bf x}|\bm{\theta})$ denote the codelength of the detail.
The total length of the MML message, $I({\bf x},\bm{\theta})$, measured in units of information (for example, bits) is the sum of the lengths of the assertion and the detail:
$I({\bf x},\bm{\theta})=\underbrace{I(\bm{\theta})}_{\rm assertion}+\underbrace{I({\bf x}|\bm{\theta})}_{\rm detail}.$ (10)
The length of the assertion measures the complexity of the model, with longer assertions able to state more parameters with high accuracy or describe more complicated model structures. In contrast, a short assertion may encode the model parameters imprecisely and describe only simple models. The length of the detail tells us how well the model stated in the assertion is able to fit (or describe) the data. A complex model with a long assertion will have lots of explanatory power and be able to encode more data strings using fewer bits compared to a simpler model. MML seeks the model
$\hat{\bm{\theta}}({\bf x})=\operatorname*{arg\,min}_{\bm{\theta}\in\bm{\Theta}}\left\{I({\bf x},\bm{\theta})\right\}$ (11)
that minimises the length of the two-part message. The key point is that minimising the two part message requires balancing the complexity of the model (assertion) with how well the model describes the data (detail). Ideally, we wish to find the simplest model that fits the observed data well enough; essentially, a formalisation of the famous razor of Occam. An advantage of MML is that the message length, measured in (say) bits, is a universal gauge that allows comparison across models with different model structures and numbers of parameters. As long as we can compute the MML codelengths of models, we can compare them. In this fashion, an MML practitioner is able to compare, for example, a linear regression model [30], to a finite mixture model [31] to a decision tree [32] via their codelengths for some observed data set.
The exact solution to (11) is known as Strict MML [33, 29], and is deemed to be the gold standard codelength. Strict minimum message length (SMML) seeks the partition $P$ of $\mathcal{X}$ that minimises the expected codelength of a two-part message describing the data ${\bf x}\in\mathcal{X}$ and a model $\bm{\theta}\in\Theta^{*}=\{\bm{\theta}_{1},\bm{\theta}_{2},\ldots\}\subset\bm{\Theta}$, with $\bm{\theta}_{j}\in\mathbb{R}^{p}$ [33, 29]. Both the parameter space $\Theta^{*}$ and the data space $\mathcal{X}$ are assumed to be countable, without loss of generality. Given a partition $P$ of the data space $\mathcal{X}$, the expected SMML codelength is
$I({\bf x},\bm{\theta})\equiv I(P)=\sum_{C\in P}f(C),$ (12)
where $f(C)$ is a expected codelength of the data ${\bf x}\in C$ in cell $C$ given by
$f(C)=\underbrace{-\sum_{{\bf x}\in C}r({\bf x})\log q(C)}_{\tt assertion}-\underbrace{\sum_{{\bf x}\in C}r({\bf x})\log p({\bf x}|\hat{\theta}(A))}_{\tt detail},\quad C\in P,$ (13)
and $r(\cdot)$ is the marginal distribution of the data
$r({\bf x})=\sum_{\theta\in\Theta}\pi(\theta)\,p({\bf x}|\theta).$ (14)
The volume of a cell, $q(C)$, is the coding probability of stating the estimate $\hat{\bm{\theta}}(C)$ for cell $C\in P$, while the estimate used for cell $C$ is obtained by minimising the expected negative log-likelihood over data ${\bf x}\in C$. Formally, we have
$q(C)=\sum_{{\bf x}\in C}r({\bf x}),\quad\hat{\bm{\theta}}(C)=\text{argmin}_{\bm{\theta}}\left\{-\sum_{{\bf x}\in C}r({\bf x})\log p({\bf x}|\theta)\right\}.$ (15)
The coding probability of the estimate for cell $C$ depends on the number of data points that are assigned to the cell, as measured by the volume $q(C)$. Specifically, the coding probability is the sum of the marginal distribution of each data point in the cell. Clearly, the larger the cell volume, the smaller the codelength for stating the estimate $\hat{\bm{\theta}}(C)$. The corresponding estimate $\hat{\bm{\theta}}(C)$ is obtained by minimising the average (with respect to the marginal distribution) negative log-likelihood of the data in the cell. The second part of the message measures how well the model $\hat{\bm{\theta}}(C)$ fits the data ${\bf x}\in C$. Observe that the second term (i.e., the detail) in $f(C)$ is the only term that depends on $\bm{\theta}$, for a given partition $P$. SMML seeks the partition $\hat{P}$ of $\mathcal{X}$ that minimises the expected codelength (12); that is,
$\hat{P}=\operatorname*{arg\,min}_{P\in\Pi^{\mathcal{X}}}I(P)$ (16)
where $\Pi^{\mathcal{X}}$ denotes the family of all partitions of the set $\mathcal{X}$. In general, to compute the optimal SMML codelength for a given sampling distribution one requires searching over all partitions $\Pi^{\mathcal{X}}$ of the data space $\mathcal{X}$. Brute force enumeration is not computationally feasible even if the data space is finite as the number of partitions of an $n$-element set $\mathcal{X}$ into exactly $k$ (non-empty) cells is the Stirling number of the second kind
$S(n,k)=\sum_{i=0}^{k}\frac{(-1)^{k-i}i^{n}}{i!(k-i)!},$ (17)
which grows rapidly for moderate values of $n$ and $k$; e.g., $S(10,5)=42,525$. Moreover, the total number of partitions of a set with $n$ elements is the $n$-th Bell number
$B_{n}=\sum_{k=0}^{n}S(n,k).$ (18)
It can be shown that $(n/4)^{n/2}\leq B_{n}\leq n^{n}$, thus Bell numbers grow exponentially with $n$ and are very large even for relatively small sets $\mathcal{X}$; (e.g., $B_{10}$ = $115,975$). Farr and Wallace [34] show that obtaining the optimal SMML codelength is, in general, an NP-hard problem.
The high computational complexity of Strict MML, renders its application, outside of simple models with a one dimensional sufficient statistic [35, 34], mostly of interest from a theoretical standpoint only. Although there exist several approximations to the Strict MML codelength, the MML87 approximation [27, 29] is perhaps the most widely applied. Under suitable regularity conditions [29]) (pp. 226), the MML87 codelength for data $\bf x$} is
$\mathcal{I}_{87}({\bf x},\bm{\theta})=\underbrace{-\log\pi(\bm{\theta})+\frac{1}{2}\log\lvert{\bf J}_{\bm{\theta}}(\bm{\theta})\rvert+\frac{P}{2}\log\kappa_{P}}_{\rm assertion}+\underbrace{\frac{P}{2}-\log p({\bf x}|\bm{\theta})}_{\rm detail}$ (19)
where $P$ is the number of free parameters, $\pi_{\bm{\theta}}(\bm{\theta})$ is the prior distribution for the parameters $\bm{\theta}$, $\lvert{\bf J}_{\bm{\theta}}(\bm{\theta})\rvert$ is the determinant of the expected Fisher information matrix, $p({\bf x}|\bm{\theta})$ is the likelihood function of the model and $\kappa_{P}$ is a quantization constant [36, 37]; for small $P$ we have
$\kappa_{1}=\frac{1}{12},\quad\kappa_{2}=\frac{5}{36\sqrt{3}},\quad\kappa_{3}=\frac{19}{192\times 2^{1/3}},$ (20)
while, for large $P$, $\kappa_{P}$ is well-approximated by [29]:
$\frac{P}{2}(\log\kappa_{P}+1)\approx-\frac{P}{2}\log 2\pi+\frac{1}{2}\log P\pi-\gamma,$ (21)
where $\gamma\approx 0.5772$ is the Euler–Mascheroni constant. Rather than searching for the partition of the data space that leads to the smallest expected codelength, a process that is known to be NP hard, MML87 approximates the coding probability of the estimate $\bm{\hat{\theta}}$ (i.e., the volume $q(C)$ of a cell $C$) as:
$q(C)\approx\pi(\hat{\bm{\theta}})\underbrace{\left(\lvert{\bf J}_{\bm{\theta}}(\hat{\bm{\theta}})\rvert\kappa_{P}^{P}\right)^{-\frac{1}{2}}}_{w(\hat{\bm{\theta})}}$, (22)
[…]
For many sufficiently well-behaved models, the MML87 codelength is virtually identical to the Strict MML codelength while being simpler to compute, requiring only the prior distribution for the model parameters and the determinant of the expected Fisher information matrix. Additionally, for large sample sizes $N\to\infty$, it is easy to show that the MML87 codelength is asymptotically equivalent to the well-known Bayesian information criterion (BIC) [38]
$\mathcal{I}_{87}({\bf x},\bm{\theta})=-\log p({\bf x}|\bm{\theta})+\frac{P}{2}\log N+O(1),$ (23)
where the $O(1)$ term depends on the prior distribution, the Fisher information and the number of parameters $p$. The MML87 codelength results in estimates that are invariant under (smooth) one-to-one reparameterisation, just like the maximum likelihood estimate. MML87 has been applied to a wide range of statistical models including decision trees [32], causal inference [39], factor analysis [23] and mixture models [31]. We next discuss how to compute the MML87 codelength approximation for the PCA model.
III-A Orthogonality constraints
As seen in Section I, it is well-known that the PCA model is not identifiable given the data. A key reason for this is that the latent vectors affect the likelihood only through their outer product ${\bf A}{\bf A}^{\prime}=\sum_{j=1}^{J}{\bf a}_{j}{\bf a}^{\prime}_{j}.$ However, there are infinitely many sets of vectors that could generate the same matrix. To resolve this ambiguity, it is a convention to estimate the factor load vectors to be mutually orthogonal; that is,
${\bf A}^{\prime}{\bf A}=\bm{\alpha}^{2}={\rm diag}(\alpha^{2}_{1},\ldots,\alpha^{2}_{J}),\quad\alpha_{j}=({\bf a}_{j}^{\prime}{\bf a}_{j})^{\frac{1}{2}},\quad(j=1,\ldots,J),$ (24)
where $\alpha_{j}$ denote the length of the $j$-th load vector. We enforce orthogonality constraints by parameterizing the matrix $\bf A$} in terms of Givens rotations [40]. Specifically, we write $\bf A$} as
${\bf A}$ $=\left[R_{12}(\phi_{1,2})\cdots R_{1,K}(\phi_{1,K})R_{2,3}(\phi_{2,3})\cdots R_{2,K}(\phi_{2,K})\cdots R_{J,J+1}(\phi_{J,J+1})\cdots R_{J,K}(\phi_{J,K}){\bf I}_{K,J}\right]\bm{\alpha}$ (25)
$={\bf R}\,\bm{\alpha},$ (26)
where ${\bf I}_{K,J}$ is the first $J$ columns of a $K\times K$ identity matrix and $R_{i,j}(\phi_{i,j})$ is a $(K\times K)$ rotation matrix that is equal to the identity matrix except for the $(i,i)$ and $(j,j)$ positions which are replaced by $\cos(\phi_{i,j})$, and the $(i,j)$ and $(j,i)$ positions which are replaced by $\sin(\phi_{i,j})$ and $\sin(\phi_{i,j})$ respectively. Thus ${\bf R}\in\mathbb{R}^{K\times J}$ and $\bm{\alpha}\in\mathbb{R}^{J\times J}$ denote the orientations and lengths of the factor load vectors, respectively. For example, when $K=J=2$, we have three free parameters
${\bf A}=\left(\begin{array}[]{cc}\cos\left(\phi_{1,2}\right)&-\sin\left(\phi_{1,2}\right)\\ \sin\left(\phi_{1,2}\right)&\cos\left(\phi_{1,2}\right)\\ \end{array}\right)\left(\begin{array}[]{cc}1&0\\ 0&1\\ \end{array}\right)\left(\begin{array}[]{c}\alpha_{1}\\ \alpha_{2}\\ \end{array}\right),$ (27)
the two factor lengths $\alpha_{1},\alpha_{2}$ and the rotation angle $\phi_{1,2}$ of the basis formed by the two factor-load directions relative to the canonical axes. This parameterisation isolates orientation and scale and explicitly takes into account that the estimated factor loads are mutually orthogonal. The model parameters are now
  • the lengths of the $J$ latent factors $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{J})\in\mathbb{R}_{+}^{J}$,
  • the orientation of the factor load vectors as captured by the D=JKJ(J+1)/2D=JK-J(J+1)/2 angles
    $\bm{\phi}=(\phi_{1,2},\ldots,\phi_{1,K},\phi_{2,3},\ldots,\phi_{2,K},\ldots,\phi_{J,J+1},\ldots,\phi_{J,K}),$
  • and the residual variance $\sigma^{2}>0$.
III-B Fisher information
Following lengthy and tedious algebra, the expected Fisher information matrix is seen to be block diagonal with determinant
$|{\bf J}(\bm{\alpha},\sigma,\bm{\phi})|$ $=N^{P}|{\bf J}(\bm{\alpha},\sigma)|\,|{\bf J}(\bm{\phi})|$ (28)
$|{\bf J}(\bm{\alpha},\sigma)|$ $=\frac{2^{J+1}(K-J)}{{\sigma^{2}}}\prod_{j=1}^{J}\frac{\alpha_{j}^{2}}{\left(\alpha_{j}^{2}+\sigma^{2}\right)^{2}}$ (29)
$|{\bf J}(\bm{\phi})|$ $=|J_{{\bf A}\to\bm{\phi}}|^{2}\left(\prod_{i=1}^{J}\left(\frac{\alpha_{i}^{4}}{\sigma^{2}}\right){}^{K-J}\frac{1}{\left(\alpha_{i}^{2}+\sigma^{2}\right){}^{K-1}}\right)\prod_{j<k}\left(\alpha_{j}^{2}-\alpha_{k}^{2}\right)^{2}$ (30)
where $|J_{{\bf A}\to\bm{\phi}}|$ is the transformation of measure under the Givens representation [40]
$|J_{{\bf A}\to\bm{\phi}}|=\prod_{i=1}^{J}\prod_{j=i+1}^{K}(\cos\phi_{i,j})^{j-i-1},$
and $P=(D+J+1)$ is the total number of free parameters. Combining all the terms we have
$|{\bf J}(\bm{\alpha},\sigma,\bm{\phi})|$ $=N^{P}\frac{\left(2^{J+1}(K-J)\right)}{{\sigma^{2}}}|J_{{\bf A}\to\bm{\phi}}|^{2}\left(\prod_{i=1}^{J}\frac{\alpha_{i}^{2}}{\left(\alpha_{i}^{2}+\sigma^{2}\right)^{2}}\left(\frac{\alpha_{i}^{4}}{\sigma^{2}}\right){}^{K-J}\frac{1}{\left(\alpha_{i}^{2}+\sigma^{2}\right){}^{K-1}}\right)\prod_{j<k}\left(\alpha_{j}^{2}-\alpha_{k}^{2}\right)^{2}$
$=\frac{N^{P}2^{J+1}(K-J)|J_{{\bf A}\to\bm{\phi}}|^{2}}{{\sigma^{2(J(K-J)+1)}}}\prod_{i=1}^{J}\frac{\alpha_{i}^{4(K-J)+2}}{\left(\alpha_{i}^{2}+\sigma^{2}\right)^{K+1}}\prod_{j<k}^{J}\left(\alpha_{j}^{2}-\alpha_{k}^{2}\right)^{2}.$ (31)
The Fisher information matrix can be singular in two specific regions of the parameter space. First, if two latent factors have identical lengths $(\alpha_{j}=\alpha_{k},j\neq k)$ the term
$\prod_{j<k}(\alpha_{j}^{2}-\alpha_{k}^{2})^{2}$
becomes zero as the two eigenvalues of the Fisher matrix become identical. This implies that the corresponding eigenvectors define a spherical subspace where any rotation within this subspace leaves the Fisher matrix unchanged. As we shall see in Section III-C, this is not a problem for the MML codelength as the problematic term, by design, cancels with a similar term in the prior distribution. The second type of singularity occurs when the MML estimate for a factor length is zero. The Fisher information contains the term
$\prod_{i=1}^{J}\alpha_{i}^{2(2(K-J)+1)}$
that leads to a vanishing determinant for any $\alpha_{j}=0$. This influences how we proceed with model selection. As we shall see in Section III-D, if the optimization drives an MML estimate $\hat{\alpha}_{j}\to 0$, we reject the model with $J$ factors and optimise for the simpler model with J1J-1 factors.
III-C Prior information
The prior distribution for the standard deviation $\sigma>0$ is chosen to be the scale-invariant density
$\pi_{\sigma}(\sigma)\propto\sigma^{-1},$ (32)
defined over some suitable range. The prior distribution for the matrix of factor loads ${\bf A}\in\mathbb{R}^{K\times J}$ is not immediately obvious as the estimates of the factor loads are enforced to be mutually orthogonal. Ideally, we would like a prior distribution that is uniform over the direction of the $J$ factors, while the distribution of the lengths of these vectors should be heavy tailed to allow for a wide range of lengths. We do not wish to make the assumption that the true factor loads are mutually orthogonal as there is no reason to believe that this would be the case a priori. Instead, we follow a similar approach to [24] and assume a prior distribution over the unknown true latent vectors that is then transformed to account for the estimated factors being mutually orthogonal. Further, as in [24], we shall consider a prior distribution for the scaled factors
${\bf b}_{j}=\left(\frac{{\bf a}_{j}}{\sigma}\right),\quad\beta_{j}=({\bf b}_{j}^{\prime}{\bf b}_{j})^{\frac{1}{2}},\qquad(j=1,\ldots,J),$
where the residual variance is used as a default scale. Let $\tilde{\bf B}\in\mathbb{R}^{K\times J}$ denote the matrix containing the $J$ true (unknown) scaled factors. We assume $\tilde{\bf B}$ to follow a matrix variate Cauchy distribution [41] with probability density function
$\pi_{\tilde{A}}(\tilde{\bf B})=\frac{\Gamma_{K}((K+J)/2)}{\pi^{KJ/2}\Gamma_{K}(K/2)}{\rm det}({\bf I}_{K}+\tilde{\bf B}\tilde{\bf B}^{\prime})^{-(K+J)/2}.$ (33)
This is a reasonable choice as the matrix variate Cauchy is spherically symmetric and has appropriately heavy tails. Further, our choice of the prior distribution implies that $\tilde{\bf B}^{\prime}\in\mathbb{R}^{J\times K}$ follows a matrix variate Cauchy distribution with density
$\pi_{\tilde{B}^{\prime}}(\tilde{\bf B}^{\prime})=\frac{\Gamma_{J}((K+J)/2)}{\pi^{KJ/2}\Gamma_{J}(J/2)}{\rm det}({\bf I}_{J}+\tilde{\bf B}^{\prime}\tilde{\bf B})^{-(K+J)/2}.$ (34)
Consider the unique matrix polar decomposition
$\tilde{\bf B}^{\prime}={\bf W}_{B}^{\frac{1}{2}}\,{\bf H}_{B},\quad{\bf W}_{B}=\tilde{\bf B}^{\prime}\tilde{\bf B},\quad{\bf H}_{B}=(\tilde{\bf B}^{\prime}\tilde{\bf B})^{-\frac{1}{2}}\tilde{\bf B}^{\prime}$, (35)
where ${\bf H}_{B}$ is defined over the Stiefel manifold $\mathcal{V}_{J}(\mathbb{R}^{K})$ and ${\bf W}_{B}$ is a symmetric positive definite matrix. We may think of the matrix ${\bf H}_{B}$ as the orientation matrix, while the matrix ${\bf W}_{B}$ determines the squared lengths of the true scaled latent vectors. If $\tilde{\bf B}^{\prime}$ follows a matrix variate Cauchy distribution, it is known that ${\bf H}_{B}$ is distributed uniformly over the Stiefel manifold with density function [41]:
$\pi_{H}({\bf H}_{B})=\frac{1}{{\rm Vol}(\mathcal{V}_{J}(\mathbb{R}^{K}))},\quad{\rm Vol}(\mathcal{V}_{J}(\mathbb{R}^{K}))=\frac{2^{J}\pi^{KJ/2}}{\Gamma_{J}(K/2)},$ (36)
where $\Gamma_{p}(y)$ is the multivariate Gamma function
$\Gamma_{J}(y)=\pi^{J(J-1)/4}\prod_{j=1}^{J}\Gamma(y+(1-j)/2).$
Further, the random variable ${\bf W}_{B}$ representing the squared lengths of the true scaled factors is independent of ${\bf H}_{B}$ with probability density function [41]
$\pi_{W}({\bf W}_{B})$ $\propto$ ${\rm det}({\bf W}_{B})^{(K-J-1)/2}{\rm det}({\bf I}_{K}+{\bf W}_{B})^{-(K+J)/2}$ (37)
which is a matrix variate beta type II distribution ${\bf W}_{B}\sim B_{J}^{II}(K/2,J/2)$ with parameters $(K/2,J/2)$ (see [42], pp. 166, for further details); this is also known as the matrix variate $F$ distribution (see, for example, [43]). Recall that the estimated (scaled) factor load vectors obey
${\bf S}_{B}=\tilde{\bf B}\tilde{\bf B}^{\prime}=\sum_{j=1}^{J}\tilde{\bm{\beta}}_{j}\tilde{\bm{\beta}}_{j}^{\prime}=\sum_{j=1}^{J}\bm{\beta}_{j}\bm{\beta}_{j}^{\prime}={\bf B}{\bf B}^{\prime},\quad\bm{\beta}_{j}^{\prime}\bm{\beta}_{k\neq j}=0.$ (38)
where ${\bf S}_{B}$ is a $(K\times K)$ symmetric matrix of rank $J$. This implies that the distribution of the squared scaled lengths $\beta_{j}^{2}$ of the estimated latent vectors is the joint distribution of the $J$ eigenvalues of ${\bf S}_{B}$ which is (see Appendix A)
$\pi_{\bm{\beta}^{2}}(\beta_{1}^{2},\ldots,\beta_{J}^{2})$ $=\frac{\pi^{J^{2}/2}}{\Gamma_{J}(J/2)\mathcal{B}_{J}(K/2,J/2)}\prod_{j=1}^{J}\beta_{j}^{(K-J-1)}(1+\beta_{j}^{2})^{-(K+J)/2}\prod_{j<k}^{J}|\beta_{j}^{2}-\beta_{k}^{2}|,$
where $\mathcal{B}_{p}(a,b)$ denote the multivariate beta function
$\mathcal{B}_{J}(a,b)=\frac{\Gamma_{J}(a)\Gamma_{J}(b)}{\Gamma_{J}(a+b)}.$
The prior distribution of the lengths of the scaled latent factors is
$\pi_{\bm{\beta}}(\beta_{1},\ldots,\beta_{J})$ $=\frac{2^{J}\pi^{J^{2}/2}}{\Gamma_{J}(J/2)\mathcal{B}_{J}(K/2,J/2)}\prod_{j=1}^{J}\beta_{j}^{(K-J)}(1+\beta_{j}^{2})^{-(K+J)/2}\prod_{j<k}^{J}|\beta_{j}^{2}-\beta_{k}^{2}|.$ (39)
Finally, the prior distribution for the lengths of the (unscaled) latent factors is
$\pi_{\bm{\alpha}}(\alpha_{1},\ldots,\alpha_{J})$ $=\frac{2^{J}\pi^{J^{2}/2}\sigma^{J^{2}}}{\Gamma_{J}(J/2)\mathcal{B}_{J}(K/2,J/2)}\prod_{j=1}^{J}\alpha_{j}^{(K-J)}(\sigma^{2}+\alpha_{j}^{2})^{-(K+J)/2}\prod_{j<k}^{J}|\alpha_{j}^{2}-\alpha_{k}^{2}|.$ (40)
The complete prior distribution over all model parameters is
$\pi(\bm{\alpha},\sigma,\bm{\phi})=\pi_{\sigma}(\sigma)\,\pi_{\bm{\alpha}}(\alpha_{1},\ldots,\alpha_{J})|J_{{\bf A}\to\bm{\phi}}|\,J!,$ (41)
where the term $J!$ is included because the labelling of the latent factors is arbitrary and $|J_{{\bf A}\to\bm{\phi}}|$ is the transformation of measure from the matrix parametrization $\bf A$} to the orthogonality-preserving parameterization based on Givens rotations.
III-D Codelength
Omitting constants, the MML codelength [27] for the probabilistic PCA model is
$\mathcal{I}$ $\propto\frac{N}{2}\log|\bm{\Sigma}|+\frac{N}{2}{\rm tr}\left(\bm{\Sigma}^{-1}{\bf S}_{x}\right)-KJ\log(\sigma)+\frac{1}{2}\sum_{j=1}^{J}\log\left[\alpha_{j}^{2(K-J+1)}\left(\alpha_{j}^{2}+\sigma^{2}\right)^{(J-1)}\right]$
where ${\bf S}_{x}=\frac{1}{N}\sum_{i}{\bf x}_{i}{\bf x}_{i}^{\prime}$ is the sample variance-covariance matrix. To obtain MML estimates, we start with the Lagrangian of the factor orientations
$\psi({\bf R})$ $=\log|\bm{\Sigma}|+{\rm tr}\left(\bm{\Sigma}^{-1}{\bf S}_{x}\right)-{\rm tr}{\bf L}({\bf R}^{\prime}{\bf R}-I),$
where $\bf L$} is a $J\times J$ symmetric matrix of Lagrange multipliers. Clearly, minimising $\psi({\bf R})$ is equivalent to minimising the codelength with respect to $\bf R$}. The first differential of the Lagrangian is
$\partial\psi({\bf R})$ $=2\text{tr}\left[\bm{\alpha}{\bf A}^{\prime}\left(\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}{\bf S}_{x}\bm{\Sigma}^{-1}\right)(d{\bf R})\right]-2\text{tr}\left({\bf L}{\bf R}^{\prime}(d{\bf R})\right),$
which implies the following first order conditions
$\bm{\alpha}{\bf A}^{\prime}\left(\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}{\bf S}_{x}\bm{\Sigma}^{-1}\right)$ $={\bf 0}$ (42)
${\bf L}{\bf R}^{\prime}$ $={\bf 0}$ (43)
${\bf R}^{\prime}{\bf R}$ $={\bf I}_{J}$ (44)
From (43) we have that ${\bf L}={\bf 0}$ and from (42)
${\bf S}_{x}{\bf R}$ $={\bf R}\,\text{diag}\left(\sigma^{2}+\alpha_{1}^{2},\ldots,\sigma^{2}+\alpha_{J}^{2}\right)$
${\bf S}_{x}{\bf r}_{j}$ $={\bf r}_{j}(\sigma^{2}+\alpha_{j}^{2}),\quad(j=1,\ldots,J).$
We see that, at the codelength minimum, the MML estimate of the factor orientations is the matrix $\bf R$} whose columns are the top $J$ eigenvectors of the variance–covariance matrix ${\bf S}_{x}$ with eigenvalues $\delta_{j}=(\sigma^{2}+\alpha_{j}^{2})$, for $j=1,\ldots J$. This is identical to the corresponding maximum likelihood estimate. Omitting constants that do not depend on the residual variance, the concentrated codelength, as a function of $\sigma^{2}$ is
$\mathcal{I}(\sigma)$ $\propto\frac{N}{2}\log\left((\sigma^{2})^{K-J}\prod_{j=1}^{J}(\alpha_{j}^{2}+\sigma^{2})\right)+\frac{N}{2\sigma^{2}}\left(\sum_{j=1}^{K}\delta_{j}\right)-\frac{N}{2\sigma^{2}}\sum_{j=1}^{J}\alpha_{j}^{2}$
$\quad-KJ\log(\sigma)+\frac{1}{2}\sum_{j=1}^{J}\log\left[\alpha_{j}^{2(K-J+1)}\left(\alpha_{j}^{2}+\sigma^{2}\right)^{(J-1)}\right]$
$=\frac{N(K-J)-KJ}{2}\log\left(\sigma^{2}\right)+\frac{N}{2\sigma^{2}}\left(\sum_{j=1}^{K}\delta_{j}\right)-\frac{N}{2\sigma^{2}}\sum_{j=1}^{J}(\delta_{j}-\sigma^{2})+\frac{(K-J+1)}{2}\sum_{j=1}^{J}\log\left(\delta_{j}-\sigma^{2}\right)$ (45)
We next discuss how to obtain the MML estimate of the residual variance from the concentrated message length.
Theorem 1.
Let $\tau=\sigma^{2}$. The concentrated codelength (45) has $(J+1)$ stationary points equal to the roots of the $n=(J+1)$-degree gradient polynomial
$P(\tau)=a_{n}\tau^{n}+a_{n-1}\tau^{n-1}+\cdots+a_{1}\tau+a_{0},\quad(0<\tau<\delta_{J})$ (46)
with coefficients
$a_{j}$ $=(-1)^{j+1}\left[\hat{\tau}_{{\rm ML}}\,e_{J-j}+\left(1-\frac{KJ-j+1}{N(K-J)}+\frac{j-1}{N}\right)e_{J-j+1}\right],\quad(0\leq j\leq J+1)$ (47)
where $\hat{\tau}_{\rm ML}$ is the maximum likelihood estimate of the residual variance and $e_{t}$ denote elementary symmetric polynomials $e_{t}(\delta_{1},\ldots,\delta_{J})$ in $J$ variables $(\delta_{1},\ldots,\delta_{J})$.
Proof.
We take the convention that $e_{t}(\cdot)=0$ for $t<0$ and $t>J$. For example, for $J=3$, we have the following four elementary symmetric polynomials
$e_{0}(\delta_{1},\delta_{2},\delta_{3})$ $=1$
$e_{1}(\delta_{1},\delta_{2},\delta_{3})$ $=\delta_{1}+\delta_{2}+\delta_{3},$
$e_{2}(\delta_{1},\delta_{2},\delta_{3})$ $=\delta_{1}\delta_{2}+\delta_{1}\delta_{3}+\delta_{2}\delta_{3},$
$e_{3}(\delta_{1},\delta_{2},\delta_{3})$ $=\delta_{1}\delta_{2}\delta_{3}.$
The concentrated codelength can be written as
$\mathcal{I}(\tau)$ $\propto\frac{N(K-J)-KJ}{2}\log\left(\tau\right)+\frac{N(K-J)\hat{\tau}_{\rm ML}}{2\tau}+\frac{(K-J+1)}{2}\sum_{j=1}^{J}\log\left(\delta_{j}-\tau\right).$
Differentiating the above with respect to $\tau$, we get
$\frac{d\mathcal{I}}{d\tau}$ $=\frac{N(K-J)-KJ}{2\tau}-\frac{N(K-J)\hat{\tau}_{\rm ML}}{2\tau^{2}}-\frac{K-J+1}{2}\sum_{j=1}^{J}\frac{1}{\delta_{j}-\tau}.$ (48)
Let A=N(KJ)KJA=N(K-J)-KJ, $B=N(K-J)\hat{\tau}_{\rm ML}$ and C=KJ+1C=K-J+1. Multiplying both sides by $2\tau^{2}$ and re-arranging:
$A\tau-B$ $=C\tau^{2}\sum_{j=1}^{J}\frac{1}{\delta_{j}-\tau}.$ (49)
Using elementary symmetric polynomials, define
$Q_{J}(\tau)=\prod_{j=1}^{J}(\delta_{j}-\tau)=\sum_{k=0}^{J}(-1)^{k}e_{J-k}\tau^{k},\qquad Q_{J}^{\prime}(\tau)=-\sum_{j=1}^{J}\prod_{k\neq j}(\delta_{k}-\tau).$ (50)
Multiplying (49) by $Q_{J}(\tau)$, we get the polynomial
$A\tau\,Q_{J}(\tau)-B\,Q_{J}(\tau)+C\tau^{2}Q_{J}^{\prime}(\tau)=0.$ (51)
The coefficients of $\tau^{m}$ for each term are
  • For $A\tau\,Q_{J}(\tau)$ with $m=1,\ldots,J+1$: $A(-1)^{m-1}e_{J-m+1}$;
  • For $B\,Q_{J}(\tau)$ with $m=0,\ldots,J$: $B(-1)^{m+1}e_{J-m}$; and
  • For $C\tau^{2}Q_{J}^{\prime}(\tau)$ with $m=2,\ldots,J+1$: $C(-1)^{m-1}(m-1)e_{J-m+1}$.
Dividing by N(KJ)N(K-J), we get the polynomial $P(\tau)$ with coefficients
$a_{m}=(-1)^{m+1}(\hat{\tau}_{\rm ML}\cdot e_{J-m}+c_{m}\cdot e_{J-m+1}),\qquad c_{m}=1-\frac{KJ-m+1}{N(K-J)}+\frac{m-1}{N},$ (52)
which matches (46) and (47). ∎
MML estimate of the residual variance $\hat{\sigma}^{2}_{\rm MML}$ is the stationary point in the interior of the parameter space $0<\tau<\delta_{J}$ that yields the shortest codelength. MML estimates of the factor lengths can be obtained from $\hat{\alpha}_{j}=(\delta_{j}-\hat{\sigma}^{2}_{\rm MML})^{\frac{1}{2}}$ for all $j=1,\ldots,J$. Since the gradient polynomial $P(\tau)$ is a continuous function of $\tau$ that is negative at $\tau=0$ and $\tau=\delta_{J}$, a root exists in the interval if and only if $P(\tau)$ has a local maximum that is strictly greater than 0 in the same interval. As will be seen in Theorem 4 and our discussion for $J=1$ below, if the signal is too weak, the real roots of this polynomial disappear or violate the condition $0<\tau<\delta_{J}$. This implies that the minimum message length solution is not found in the interior of the $J$-factor model parameter space and is instead found in the J1J-1 model space (i.e., a model with one less latent factor). The next theorem characterises the roots of the gradient polynomial.
Theorem 2.
Let $\mathcal{I}(\tau)$ denote the concentrated codelength (45) defined on the domain $(0,\delta_{J})$ and let
$h(\tau)=2\tau^{2}\left(\frac{d\mathcal{I}(\tau)}{d\tau}\right)=L(\tau)-R(\tau),\quad L(\tau)=A\tau-B,\quad R(\tau)=C\tau^{2}\sum_{j=1}^{J}(\delta_{j}-\tau)^{-1},$ (53)
where $A=N(K-J)-KJ,B=N(K-J)\hat{\tau}_{\rm ML}$ and C=KJ+1C=K-J+1. The solution space of the MML estimate exhibits a phase-transition behavior that can be classified into two regimes:
  1. 1.
    Weak signal ($\delta_{J}\approx\hat{\tau}_{\rm ML}$): the codelength minimum occurs at the boundary $\delta_{J}$, which implies zero real roots in the domain $(0,\delta_{J})$. The model with $J$ factors is rejected in favour of the simpler model with J1J-1 factors; and
  2. 2.
    Strong signal ($\delta_{J}\gg\hat{\tau}_{\rm ML}$): the codelength exhibits two stationary points in $(0,\delta_{J})$: (i) a local minimum of the codelength, which is the valid MML estimate near $\hat{\tau}_{\rm ML}$, and (ii) a local maximum of the codelength located near the singularity $\delta_{J}$.
Proof.
Consider the intersection $L(\tau)=R(\tau)$ of the linear function $L(\tau)$ with the rational function $R(\tau)$ at the boundary of the domain $(0,\delta_{J})$. The two functions do not intersect at $\tau=0$ since $L(0)<0$ and $R(0)=0$. Conversely, at the boundary $\tau=\delta_{J}$, the linear function $L(\delta_{J})$ is finite, while $R(\delta_{J})\to+\infty$. Since $R(\tau)$ is strictly convex, there are either zero intersections $(L(\tau)<R(\tau))$ or exactly two intersections in the domain $(0,\delta_{J})$. In the case of weak signal, we have $\hat{\tau}_{\rm ML}\approx\delta_{J}$, so that $h(\tau)<0$ everywhere in the domain $(0,\delta_{J})$. This implies that $\mathcal{I}(\tau)$ is strictly monotonically decreasing with the minimum occurring at $\tau\to\delta_{J}$, resulting in no solutions and a collapse of the $J$-factor model. In the case of strong signal, assume that $\delta_{J}\gg\hat{\tau}_{\rm ML}$. At the midpoint $\tau^{*}=\delta_{J}/2$, we have
$L(\tau^{*})$ $=(N(K-J)-KJ)\left(\frac{\delta_{J}}{2}\right)-N(K-J)\hat{\tau}_{\rm ML}=N(K-J)\left(\frac{\delta_{J}}{2}-\hat{\tau}_{\rm ML}\right)+O(1),$
$R(\tau^{*})$ $=(K-J+1)\left(\frac{\delta_{J}}{2}\right)^{2}\sum_{j=1}^{J}\frac{1}{\delta_{J}-\delta_{J}/2}=O(1).$
For large $N$, $L(\tau^{*})\gg R(\tau^{*})$ and so $h(\tau^{*})>0$. Since $h(0)<0$ and $h(\tau^{*})>0$, there is at least one root $\tau_{1}$ in $(0,\tau^{*})$ that is a local minimum. This is our MML estimate of the residual variance. Similarly, since $h(\tau^{*})>0$ and $h(\delta_{J})\to$-$\infty$, there is at least one root $\tau_{2}$ in $(\tau^{*},\delta_{J})$ that is a local maximum. Because $L(\tau)$ is linear and $R(\tau)$ is strictly convex, there are exactly two roots in $(0,\delta_{J})$. Re-arranging $L(\tau^{*})\gg R(\tau^{*})$ for $N$, we observe how the sample size scales with the signal to noise ratio $\rho_{J}=\delta_{J}/\hat{\tau}_{\rm ML}$:
$N\gg\frac{\delta_{J}}{\delta_{J}-2\hat{\tau}_{\rm ML}}=\frac{\rho_{J}}{\rho_{J}-2}.$
For strong signal, the right hand side approaches $N\gg 1$, while for weak signal $N\to\infty$. ∎
In the limit as $N\to\infty$ the gradient polynomial $P(\tau)$ can be factored as follows
$P(\tau)=(\tau-\hat{\tau}_{\rm ML})\prod_{j=1}^{J}(\tau-\delta_{j}).$ (54)
The $(J+1)$ roots of $P(\tau)$ are the $J$ largest eigenvalues of the sample covariance matrix and the maximum likelihood estimator of the residual variance. As the codelength is only defined when $0<\tau<\delta_{J}$, we see that, in the limit as $N\to\infty$, the minimum message length estimate of the residual variance is equal to the maximum likelihood estimate, as expected. The next theorem discusses the bias of the MML estimate of the residual variance.
Theorem 3.
Let $\rho_{j}=\alpha_{j}^{2}/\sigma^{2}$ denote the signal-to-noise ratio for the $j$-th factor in the PCA model with $J$ true latent factors. Assuming fixed $K,J$ and $N\to\infty$, the bias of the MML estimate of residual variance $\tau:=\sigma^{2}$ is:
$\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}=\frac{\tau}{N(K-J)}\left(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2}).$ (55)
Proof.
The derivation uses first-order perturbation theory around the maximum likelihood root. All expectations are taken under fixed $K$, $J$ and $N\to\infty$, and higher-order covariance terms between the sample eigenvalues and $\hat{\tau}_{\rm ML}$ are $O(N^{-2})$ and therefore neglected. The MML estimate of $\tau$ is a stationary point of the polynomial
$P(\tau)=\sum_{j=1}^{J}a_{j}\tau^{j}=0,\qquad a_{j}=(-1)^{j+1}\left(\hat{\tau}_{\rm ML}e_{J-j}+c_{j}e_{J-j+1}\right),$
where
$c_{j}=1+\epsilon_{j},\quad\epsilon_{j}=\frac{j-1}{N}-\frac{KJ-j+1}{N(K-J)}=\frac{(j-1)(K-J+1)-KJ}{N(K-J)}.$
The coefficients of this polynomial converge such that a root is exactly $\hat{\tau}_{ML}$ for $N\to\infty$. For finite $N$, we will approximate the MML estimate of $\tau$ as the MLE estimate plus a small correction term $\Delta$. The bias of the maximum likelihood estimate of $\tau$ up to second order [44] is
$\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}=-\frac{\tau}{N}\sum_{j=1}^{J}\left(1+\rho_{j}^{-1}\right)+O(N^{-2}).$ (56)
Let $A(\tau)$ denote the characteristic polynomial of the $J$ sample eigenvalues
$A(\tau)=\prod_{j=1}^{J}(\tau-\delta_{j})=\sum_{k=1}^{J}(-1)^{k}e_{J-k}\tau^{k}.$ (57)
Expand $P(\tau)$ around the MLE estimate $\hat{\tau}_{\rm ML}$
$P(\tau+\Delta)\approx P(\hat{\tau}_{\rm ML})+P^{\prime}(\hat{\tau}_{\rm ML})\Delta=0,$
where $\Delta$ is a small perturbation of order $O(1/N)$. Solving for $\Delta$, we obtain
$\Delta=-\frac{P(\hat{\tau}_{\rm ML})}{P^{\prime}(\hat{\tau}_{\rm ML})}.$ (58)
Next, we simplify the numerator and denominator using the properties of symmetric polynomials to express $\Delta$ in terms of the sample eigenvalues. As $N\to\infty$, the perturbation terms $\epsilon_{j}\to 0$, and the asymptotic form of the coefficients is
$\bar{a}_{j}=(-1)^{j+1}\left(\hat{\tau}_{\rm ML}e_{J-j}+e_{J-j+1}\right)$
Noting that
$\sum_{j=1}^{J}(-1)^{j+1}\hat{\tau}_{\rm ML}e_{J-j}\tau^{j}$ $=-\hat{\tau}_{\rm ML}\sum_{j=1}^{J}(-1)^{j}e_{J-j}\tau^{j}=-\hat{\tau}_{\rm ML}A(\tau)$
$\sum_{j=1}^{J}(-1)^{j+1}e_{J-j+1}\tau^{j}$ $=\sum_{k=0}^{J}(-1)^{k+2}e_{J-k}\tau^{k}=\sum_{k=0}^{J}(-1)^{k+2}e_{J-k}\tau^{k+1}=\tau\sum_{k=0}^{J}(-1)^{k}e_{J-k}\tau^{k},=\tau A(\tau)$
the limiting polynomial and can be written as
$\bar{P}(\tau)=(\tau-\hat{\tau}_{\rm ML})A(\tau).$
Differentiating with respect to $\tau$
$\bar{P}^{\prime}(\tau)=A(\tau)+(\tau-\hat{\tau}_{\rm ML})A^{\prime}(\tau).$ (59)
and evaluating the derivative at $\hat{\tau}_{\rm ML}$, we get
$\bar{P}^{\prime}(\hat{\tau}_{\rm ML})=A(\hat{\tau}_{\rm ML}).$
The original polynomial, evaluated at $\hat{\tau}_{\rm ML}$ is
$P(\hat{\tau}_{\rm ML})=\bar{P}(\hat{\tau}_{\rm ML})+Q(\hat{\tau}_{\rm ML})=Q(\hat{\tau}_{\rm ML}),\qquad Q(\tau)=\sum_{j=1}^{J}(-1)^{j+1}\epsilon_{j}e_{J-j+1}\tau^{j}.$
Let k=j1k=j-1 and write the polynomial $Q(\tau)$ as
$Q(\tau)$ $=\sum_{k=0}^{J}(-1)^{k+2}\left(\frac{k(K-J+1)-KJ}{N(K-J)}\right)e_{J-k}\tau^{k+1}$
$=\frac{\tau}{N(K-J)}\sum_{k=0}^{J}(-1)^{k}\left(k(K-J+1)-KJ\right)e_{J-k}\tau^{k}$
$=\frac{\tau}{N(K-J)}\left((K-J+1)\tau A^{\prime}(\tau)-KJA(\tau)\right).$
Substituting the above into our equation for $\Delta$, we get
$\Delta$ $=-\frac{\frac{\tau_{\rm ML}}{N(K-J)}\left((K-J+1)\tau A^{\prime}(\tau_{\rm ML})-KJA(\tau_{\rm ML})\right)}{A(\hat{\tau}_{\rm ML})}$
$=\frac{\hat{\tau}_{\rm ML}}{N(K-J)}\left(KJ-(K-J+1)\sum_{j=1}^{J}\frac{\hat{\tau}_{\rm ML}}{\hat{\tau}_{\rm ML}-\delta_{j}}\right).$
Next, replace sample quantities with their population counterparts and compute the expectation
$\mathbb{E}\{\Delta\}$ $=\frac{\tau}{N(K-J)}\left(KJ+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}\right).$
Recall that the MML estimate of $\tau$ is modelled as the MLE plus a small correction. We finally combine expectations and simplify to obtain the bias estimate up to second order:
$\mathbb{E}\{\hat{\tau}_{\rm MML}\}$ $=\mathbb{E}\{\hat{\tau}_{\rm MLE}\}+\mathbb{E}\{\Delta\}$
$=\tau-\frac{\tau}{N}\sum_{j=1}^{J}\left(1+\rho_{j}^{-1}\right)+\frac{\tau}{N(K-J)}\left(KJ+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2})$
$=\tau+\frac{\tau}{N(K-J)}\left(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2}).$
The maximum likelihood estimate is negatively biased and underestimates noise, with bias of order $O(1/N)$ and approximately proportional to $J\tau/N$. In contrast, the MML estimate of $\tau$ is positively biased and overestimates noise in finite samples, with bias of order $O(1/(NK))$ and approximately proportional to $J^{2}\tau/(NK)$. The absolute ratio $R$ of the two biases is:
$R=\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-J)\left(\frac{J+\sum_{j=1}^{J}\rho_{j}^{-1}}{J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}}\right).$ (60)
Observe that the MML estimate reduces bias compared to the maximum likelihood estimate by a factor roughly proportional to the dimension $K$. If the signals are strong (high signal-to-noise ratio) with $\rho_{j}\to\infty$ ($j=1,\ldots,J)$, the ratio is approximately $R\approx(K-J)/J$, suggesting that the MML estimate reduces bias by a factor proportional to the ratio of the total dimension to the latent dimension. In contrast, for weak signals with $\rho_{j}\to 0$, the terms $\rho_{j}^{-1}$ dominate and the bias reduction factor for the MML estimate is $R\approx(K-J)$. Thus, even in the case of weak signals, where the bias of both MML and maximum likelihood estimate is inflated, the relative advantage of the MML estimate remains constant. Next, we analyse model selection properties of the MML codelength.
Theorem 4.
Consider the PCA model with $J$ true latent factors. Assume that $K,J$ are fixed and that $N\to\infty$. The MML estimator detects the $j$-th latent factor (i.e., estimates a non-zero signal strength ($\hat{\alpha}_{j}>0$) if and only if the $j$-th sample eigenvalue $\delta_{j}$ exceeds a specific critical threshold relative to the estimated residual variance $\hat{\tau}_{\rm MML}$ that is given by
$\delta_{j}>\hat{\tau}_{\rm MML}\left(1+\sqrt{\frac{K_{j}}{N}}\right)^{2}+O(N^{-1}),$ (61)
where $K_{j}=K-j+1$ denotes the effective degrees of freedom available for the $j$-th eigenvector.
Proof.
The prior density for the $j$-th factor is proportional to
$\pi(\alpha_{j})\propto\alpha_{j}^{K-J}\lambda_{j}^{-(K+J)/2}.$
The determinant of the Fisher information matrix contribution for the $j$-th factor is
$|J|\propto\frac{\alpha_{j}^{2((K-J)+1)}}{\lambda_{j}^{K+1}}.$
Combining and simplifying, we get the approximate cost of coding the $j$-th factor
$-\log\alpha_{j}^{K-J}\lambda_{j}^{-(K+J)/2}+\frac{1}{2}\log\frac{\alpha_{j}^{2((K-J)+1)}}{\lambda_{j}^{K+1}}$ $=(K-J+1)\log\alpha_{j}+\frac{J-1}{2}\log\lambda_{j}$ (62)
$=\frac{K-J+1}{2}\log(\lambda_{j}-\tau)+O(1),$ (63)
where the contribution of the term $(J-1)/2\log\lambda_{j}$ can be viewed as $O(1)$ and is ignored. Let $K_{j}=(K-j+1)$ denote the effective degrees of freedom available to the $j$-th eigenvector. From the Gaussian likelihood of the PCA model, the data coding cost for the $j$-th component is:
$\frac{N}{2}\left(\frac{\delta_{j}}{\lambda_{j}}+\log\lambda_{j}\right).$
Combining the data and cost of coding a factor, we get
$\frac{N}{2}\left(\frac{\delta_{j}}{\lambda_{j}}+\log\lambda_{j}\right)+\frac{K_{j}}{2}\log(\lambda_{j}-\tau).$
Differentiating the expression with respect to $\lambda_{j}$ and re-arranging
$\frac{N}{2}\left(\frac{1}{\lambda_{j}}-\frac{\delta_{j}}{\lambda_{j}^{2}}\right)+\frac{K_{j}}{2(\lambda_{j}-\tau)}$ $=0$
$\delta_{j}$ $=\lambda_{j}+\frac{K_{j}}{N}\frac{\lambda_{j}^{2}}{\lambda_{j}-\tau}.$
At the detection threshold, the signal strength may be assumed to be small so that $\lambda_{j}=\tau(1+\gamma)$, where $\gamma>0$ is a small signal-to-noise ratio. Substituting this into the previous expression and re-arranging we find
$\frac{\delta_{j}}{\tau}=(1+\gamma)+\frac{K_{j}}{N}\frac{(1+\gamma)^{2}}{\gamma}$
Differentiating the RHS and solving for the critical point $\gamma_{*}$
$\gamma^{2}_{*}=\frac{\epsilon}{1+\epsilon}=\epsilon+O(\epsilon^{2}),\qquad\epsilon=\frac{K_{j}}{N}.$
Substituting the critical point back into the RHS, we have
$\frac{\delta_{j}}{\tau}=(1+\sqrt{\epsilon})+\epsilon\frac{(1+\sqrt{\epsilon})^{2}}{\sqrt{\epsilon}}=\left(\sqrt{\epsilon}+1\right)\left(\epsilon+\sqrt{\epsilon}+1\right)\approx\left(\sqrt{\epsilon}+1\right)^{2}$ (64)
Substituting and simplifying
$\delta_{j}>\tau\left(1+\sqrt{\frac{K_{j}}{N}}\right)^{2}+O(N^{-1}),$
concludes the proof. ∎
Observe that the MML threshold has the same functional form as the well-known Baik-Ben Arous-Péché (BBP) [25] phase transition point
$\lambda_{\rm edge}=\tau(1+\sqrt{\gamma})^{2},\qquad\gamma=K/N,$ (65)
from random matrix theory, which is derived assuming both $N,K\to\infty$ with $K/N\to\gamma$. The MML estimator tends to prune any factor that is spectrally indistinguishable from the bulk noise, while retaining factor that statistically protrude from the noise bulk.
Single latent factor. For the PCA model with a single true latent factor ($J=1$), the stationary points of the concentrated codelength are the roots of the quadratic polynomial in $\tau$:
$-\delta_{1}\hat{\tau}_{\text{ML}}+\left(\hat{\tau}_{\text{ML}}+c\delta_{1}\right)\tau-\tau^{2}=0,\quad c=1-\frac{K}{N(K-1)},$ (66)
given by
$\frac{1}{2}\left(\hat{\tau}_{\text{ML}}+c\,\delta_{1}\pm\Delta^{\frac{1}{2}}\right),\quad\Delta=c^{2}\delta_{1}^{2}+2(c-2)\delta_{1}\hat{\tau}_{\text{ML}}+\hat{\tau}_{\text{ML}}^{2}.$
where the ML estimate of $\tau$ is given in (9). The quadratic polynomial has two positive real roots if
$\frac{\delta_{1}}{\hat{\tau}_{\text{ML}}}>\left(\sqrt{\frac{K}{(K-1)N}}-1\right)^{-2}=(1+\sqrt{\gamma_{\rm MML}})^{2},\qquad\gamma_{\rm MML}=\left(\frac{\sqrt{1-c}+1-c}{c}\right)^{2}.$ (67)
Note that MML requires the ratio of the top eigenvalue to the residual variance estimate to be greater than a constant, which depends on $N$ and $K$ only, for the single-factor model to be estimable. This condition ensures separation of the largest eigenvalue from the remainder of the ‘noisy’ (i.e., bulk) eigenvalues. For example, when $N=25$ and $K=4$, the quadratic will have two real roots if
$\frac{\delta_{1}}{\hat{\tau}_{\text{ML}}}>\frac{75}{79-20\sqrt{3}}\approx 1.691.$
In the limit as $N\to\infty$, the two roots of the quadratic are $\hat{\tau}_{\rm ML}$ and $\delta_{1}$, which shows that the MML estimate of the residual noise converges to the maximum likelihood estimate, as expected. The absolute bias ratio of the maximum likelihood estimate of $\tau$ to the MML estimate simplifies to
$\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-1)\left(\frac{1+\sum_{j=1}^{J}\rho_{j}^{-1}}{1^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}}\right)=K-1.$ (68)
For one true latent factor, the MML estimate reduces bias by the factor of (K1)(K-1) irrespective of the signal-to-noise ratio. When the signal is very weak, both biases blow up in magnitude because the spike is barely separated from the noise bulk. The maximum likelihood estimate severely underestimates $\tau$, while MML slightly overestimates it, but by a factor 1/(K1)1/(K-1). When the signal is very strong, both biases shrink to $O(1/N)$ constants with the maximum likelihood bias being negative and about (K1)(K-1) times larger than the MML bias.
In the asymptotic regime where $N\to\infty$ with known residual noise $\sigma^{2}$, the likelihood ratio test for the single factor model depends only on the largest sample eigenvalue $\delta_{1}$ [45, 46]. The MML procedure may be seen as equivalent to the generalised likelihood ratio test (GLRT), specifically the largest root test with residual noise $\hat{\sigma}^{2}$ estimated, rather than known. Bianchi et al. [47] develop a GLRT based on the test statistic $T_{N}=\frac{\delta_{1}}{\frac{1}{K}\sum_{j}\delta_{j}}$ (see their Proposition (1)) where the null hypothesis (i.e., the no-factor model) is rejected for large values of $T_{N}$; note that $T_{N}$ is equivalent, up to a non-linear monotonic transformation, to $\delta_{1}/\hat{\tau}_{\rm ML}$ [46].
In this limiting case, the sample eigenvector associated with the largest sample eigenvalue is a consistent estimate of the corresponding population eigenvector only if $K/N\to 0$ [48]. The asymptotic joint distribution of $\delta_{1}/\hat{\tau}_{\rm ML}$ is derived in [49] who also use this to construct a sequence of hypothesis tests for estimating the number of principal components.
Two latent factors. For a PCA model with two latent factors ($J=2$), the stationary points of the concentrated codelength are the roots of the cubic polynomial in $\tau$:
$-\delta_{1}\delta_{2}\hat{\tau}_{\text{ML}}+\left((\delta_{1}+\delta_{2})\hat{\tau}_{\text{ML}}+c_{1}\delta_{1}\delta_{2}\right)\tau-\left(\hat{\tau}_{\text{ML}}+\left(c_{0}+c_{1}\right)(\delta_{1}+\delta_{2})\right)\tau^{2}+(2c_{0}+c_{1})\tau^{3}$
where the constants
$c_{0}=\frac{K-1}{N(K-2)},\quad c_{1}=1-\frac{2K}{N(K-2)},$
depend on $N$ and $K$ only and the ML estimate of $\tau$ is given in (9). Recall that the bias of the MML estimate of $\tau$ is approximately proportional to $J^{2}\tau/(NK)$. The term $J^{2}$ implies that adding a second factor is strictly more expensive, with the estimator requiring stronger evidence to ‘upgrade’ a model from $J=1$ to $J=2$, than the upgrade from $J=0$ to $J=1$.
IV Experiments
IV-A Parameter estimation
This section compares the newly derived MML parameter estimates for the probabilistic PCA model to the standard approach based on the maximum likelihood estimator. Since MML and maximum likelihood estimates of the factor lengths (for a given $\sigma^{2}$) and factor orientations are identical, the key difference between to two approaches is in the estimation of the residual variance. Our simulation experiments are loosely based on Section 6 in [24]. We conducted $10^{5}$ simulations for each combination of the sample size $N\in\{25,50,100\}$, the number of estimated latent factors $J\in\{1,2,4\}$ and the average signal-to-noise ratio (SNR)
$\textrm{SNR}=\frac{1}{K\sigma^{2}}\sum_{j=1}^{J}\alpha_{j}^{2},$ (69)
where the dimensionality of the data was fixed to $K=10$ for all experiments. The factor directions were randomly sampled from a unit $K$-sphere while the factor lengths were randomly sampled from a half-Cauchy distribution ensuring a wide range of generating models.
NN SNR $J$ $S_{1}$ $S_{2}$ KL Divergence
MLE MML87 MLE MML87 MLE MML87
25 0.5 1 -0.027 0.000 0.003 0.002 0.246 0.225
2 -0.092 -0.013 0.011 0.003 0.486 0.356
4 -0.151 0.003 0.029 0.004 0.739 0.402
1 1 -0.025 0.000 0.003 0.002 0.248 0.231
2 -0.082 -0.008 0.010 0.003 0.492 0.378
4 -0.144 0.012 0.027 0.005 0.789 0.451
4 1 -0.023 0.000 0.003 0.002 0.253 0.237
2 -0.068 -0.002 0.008 0.003 0.500 0.407
4 -0.134 0.030 0.024 0.006 0.892 0.550
8 1 -0.023 0.000 0.003 0.002 0.254 0.239
2 -0.063 -0.000 0.007 0.003 0.503 0.418
4 -0.129 0.038 0.023 0.007 0.933 0.597
50 0.5 1 -0.013 0.000 0.001 0.001 0.116 0.111
2 -0.051 -0.010 0.004 0.002 0.230 0.192
4 -0.111 -0.004 0.015 0.002 0.385 0.238
1 1 -0.012 0.000 0.001 0.001 0.117 0.113
2 -0.045 -0.007 0.004 0.002 0.228 0.196
4 -0.105 0.002 0.014 0.002 0.398 0.258
4 1 -0.012 0.000 0.001 0.001 0.118 0.114
2 -0.036 -0.003 0.003 0.001 0.227 0.203
4 -0.093 0.011 0.011 0.002 0.418 0.295
8 1 -0.011 -0.000 0.001 0.001 0.118 0.115
2 -0.033 -0.001 0.003 0.001 0.227 0.205
4 -0.086 0.014 0.010 0.003 0.424 0.312
100 0.5 1 -0.007 0.000 0.001 0.001 0.056 0.055
2 -0.029 -0.007 0.002 0.001 0.112 0.100
4 -0.077 -0.007 0.007 0.001 0.196 0.136
1 1 -0.006 0.000 0.001 0.001 0.057 0.056
2 -0.025 -0.005 0.001 0.001 0.111 0.101
4 -0.071 -0.005 0.006 0.001 0.198 0.144
4 1 -0.006 0.000 0.001 0.001 0.057 0.056
2 -0.020 -0.002 0.001 0.001 0.109 0.102
4 -0.058 0.001 0.005 0.001 0.200 0.158
8 1 -0.006 0.000 0.001 0.001 0.057 0.056
2 -0.018 -0.001 0.001 0.001 0.109 0.102
4 -0.052 0.003 0.004 0.001 0.199 0.163
TABLE I: Performance metrics for maximum likelihood (MLE) and MML87 estimates of residual variance $\sigma^{2}$ computed over $10^{5}$ simulations.
We used the three performance metrics to evaluate the estimators:
$S_{1}=\log\left(\frac{\hat{\sigma}_{i}}{\sigma_{i}}\right),\quad S_{2}=\left(\log\frac{\hat{\sigma}_{i}}{\sigma_{i}}\right)^{2},\quad(i=1,\ldots,10^{5}),$
and the Kullback–Leibler (KL) divergence [50] between two zero-mean multivariate Gaussian distributions
$\text{KL}(\bm{\Sigma}_{0},\bm{\Sigma}_{1})=\frac{1}{2}\left(\text{tr}\left(\bm{\Sigma}^{-1}_{1}\bm{\Sigma}_{0}\right)+\log\left(\frac{\bm{|\Sigma}_{1}|}{|\bm{\Sigma}_{0}|}\right)-K\right),$
which only depends on the variance-covariance matrices of the two models. The first metric $S_{1}$ is a measure of bias, while $S_{2}$ measures estimation error in any direction. Both $S_{1}$ and $S_{2}$ are zero for exact estimates. The error measures were specifically chosen as they do not depend on the number of estimated latent vectors $J$. Simulation results averaged over $10^{5}$ iterations are shown in Table I.
The MML estimate of the residual variance was found to be superior to the usual maximum likelihood estimate for all tested combinations of sample sizes, data dimensionality and the number of latent vectors. Maximum likelihood underestimated the residual variance more strongly compared to the minimum message length estimate. The differences in the performances of the two estimates were most pronounced when the sample size was small, with a high signal-to-noise ratio (SNR) ($N\leq 50$, SNR > 4). This agrees with our earlier analysis and the theoretical findings by Kritchman and Nadler [16] who show that the maximum likelihood estimate is biased downward; even in the case of the single-factor model, the bias is significant with small sample size $N$ and remains when the SNR is large. In comparison, the MML estimate exhibits significantly less bias, even for small sample sizes.
IV-B Model selection
Next we compared the performance of MML model selection against the popular Bayesian information criterion (BIC), Laplace’s method for approximating the marginal distribution of the data [51], referred to as ‘Bayes’ henceforth, parallel analysis algorithm (SPA) [15] and the generalized information criterion (GIC) [14]. Bayes and BIC were included as popular Bayesian model selection criteria, while SPA is a permutation-based approach rooted in random matrix theory and GIC is an improved variant of the Akaike information criterion [52] for the PCA model. Using numerical experiments, [51] demonstrated that approximating Bayesian evidence is superior to methods like cross validation.
The simulation setup was identical to Section IV-A except the sample size was $N\in\{50,100\}$, the dimensionality of the data $K=10$ and the number of estimated latent factors $J\in\{1,2,4\}$. Each criterion was asked to select the best model among candidates which had between $1$ and $5$ latent factors. Along with the three performance metrics discussed in Section IV-A we also recorded how often each criteria correctly estimated the true number of latent factors. Simulation results, averaged over $10^{5}$ iterations, are shown in Table II (for SNR=1) and in Table III (for SNR = 8). Both MML and the Bayes method have similar performance and both improve significantly over the popular BIC criterion. Importantly, even when SPA or GIC select the true model with a higher proportion compared to MML, the corresponding KL divergence of the MML criterion is often lower, suggesting that the MML model is superior.
NN $J$ Method KL Divergence Model Selection (%)
<J<J =J=J >J>J
50 1 MML 0.116 97.84 2.16
BIC 0.117 99.96 0.04
Bayes 0.126 95.64 4.37
SPA 0.117 100.00 0.00
GIC 0.128 95.41 4.59
2 MML 0.178 62.46 28.59 8.95
BIC 0.190 73.01 26.97 0.02
Bayes 0.187 59.31 38.43 2.26
SPA 0.233 86.84 13.16 0.00
GIC 0.194 58.88 36.30 4.82
4 MML 0.225 77.60 20.39 2.01
BIC 0.261 99.94 0.06 0.00
Bayes 0.247 98.23 1.47 0.30
SPA 0.304 100.00 0.00 0.00
GIC 0.258 94.24 3.99 1.76
100 1 MML 0.057 99.10 0.90
BIC 0.057 100.00 0.00
Bayes 0.060 97.06 2.94
SPA 0.057 100.00 0.00
GIC 0.062 95.66 4.35
2 MML 0.088 56.07 40.36 3.57
BIC 0.095 64.78 35.22 0.00
Bayes 0.091 52.41 45.78 1.80
SPA 0.144 82.92 17.08 0.00
GIC 0.094 50.86 44.57 4.56
4 MML 0.120 85.28 4.37 10.35
BIC 0.139 99.77 0.23 0.00
Bayes 0.126 96.80 2.91 0.29
SPA 0.201 99.99 0.01 0.00
GIC 0.130 92.30 5.88 1.82
TABLE II: Model selection simulation results for minimum message length (MML), Laplace’s method for estimating Bayesian evidence, signflip parallel analysis (SPA) and a generalized information criterion (GIC) averaged over $10^{5}$ simulations. In all experiments, data dimensionality was $K=10$. SNR = 1
NN $J$ Method KL Divergence Model Selection (%)
<J<J =J=J >J>J
50 1 MML 0.119 97.58 2.42
BIC 0.118 99.96 0.04
Bayes 0.128 95.66 4.34
SPA 0.118 100.00 0.00
GIC 0.129 95.72 4.28
2 MML 0.216 30.82 44.15 25.03
BIC 0.208 36.26 63.68 0.06
Bayes 0.212 29.50 67.18 3.32
SPA 1.041 83.74 16.26 0.00
GIC 0.218 29.48 65.04 5.48
4 MML 0.299 33.69 51.57 14.75
BIC 0.342 86.68 13.25 0.07
Bayes 0.335 78.92 19.36 1.72
SPA 1.519 99.98 0.02 0.00
GIC 0.346 73.43 21.68 4.89
100 1 MML 0.057 99.03 0.97
BIC 0.057 100.00 0.00
Bayes 0.060 97.03 2.97
SPA 0.057 100.00 0.00
GIC 0.062 95.83 4.17
2 MML 0.101 27.56 63.64 8.80
BIC 0.101 32.31 67.69 0.00
Bayes 0.101 25.78 71.70 2.51
SPA 0.853 80.79 19.20 0.00
GIC 0.104 25.10 69.67 5.24
4 MML 0.156 41.40 19.37 39.23
BIC 0.168 81.85 18.14 0.01
Bayes 0.160 72.96 25.62 1.41
SPA 1.302 99.96 0.04 0.00
GIC 0.164 67.92 27.57 4.51
TABLE III: Model selection simulation results for minimum message length (MML), Laplace’s method for estimating Bayesian evidence, signflip parallel analysis (SPA) and a generalized information criterion (GIC) averaged over $10^{5}$ simulations. In all experiments, data dimensionality was $K=10$. SNR = 8
V Discussion
This manuscript derives the minimum message length (MML) codelength for the multivariate Gaussian probabilistic principal component analysis (PCA) model [6]. Although the MML estimates of the factor orientations are identical to the usual maximum likelihood (ML) estimates, an important difference between the two approaches is in the estimation of the residual variance. In this respect, minimisation of the MML codelength has two key advantages for the practitioner: (1) automatic selection of the number of principal components; and (2) an improved estimate of the residual variance. The experiments in Section IV-A demonstrated that the MML estimate of the residual variance improves upon the usual maximum likelihood estimate in terms of bias, squared error and Kullback–Leibler divergence. Unlike the MML estimate of residual variance, the maximum likelihood estimate tends to severely underestimate the residual variance [16]. These improvements over the ML estimate are substantial when the sample size is small relative to the number of parameters in the model (see Table I).
As noted above, minimising the codelength also allows automatic selection of the number of principal components in the model. We observe that model selection guided by the MML codelength is more accurate than the popular Bayesian information criterion (BIC) approach and is at least as good as the Laplace approximation to the Bayesian posterior distribution [51]. Table II shows that the MML gains in model selection accuracy over BIC are substantial in small to moderate sample sizes. As expected, as the sample size gets larger, MML codelength reduces to BIC, which is known to be consistent in large $N$ problems with a fixed number of parameters. Importantly, using the codelength to discriminate between competing hypotheses provides another advantage over BIC. Unlike BIC, MML considers the complexity of the model via the assertion part of the message and does not simply use a count of the model parameters as a surrogate for model complexity.
Further, we believe that our choice of the prior distribution over the factor load matrix is preferred to the standard Bayesian approach of assuming that the true latent factors are mutually orthogonal. There appears to be no reason to suspect that the true latent vectors are mutually orthogonal and we instead advocate for a rotation‑invariant, heavy‑tailed distribution, such as the matrix variate Cauchy distribution.
While any reasonable Bayesian approach to the PCA model with sensible priors is expected to yield similar performance to our MML codelength, MML also provides the practitioner with point estimates for all model parameters. Unlike the maximum a posteriori estimate, the MML estimates are invariant to reparameterisation of the model and are obtained by minimising the codelength. Importantly, the MML codelength is a universal yardstick that allows comparison of models across different model structures (e.g., generalized linear model [53] vs a decision tree [32]) and numbers of parameters. This means that we can use the MML codelength to discriminate between multivariate Gaussian models with specific covariance structures. For example, we can use the MML codelength to test the hypothesis that the covariance matrix is spherical versus a more general covariance structure (e.g., the PCA model).
Additionally, the MML codelength derived in this manuscript allows the PCA model to be incorporated into other component-based models with all the advantages of MML (i.e., automatic model selection and improved parameter estimation). For example, we could use the MML PCA codelength in the leaves of a decision tree, similar to the Max-Cut model in [54] or within finite mixture models of probabilistic principal component analyzers, similar to [55, 56, 57].
A. Joint eigenvalue distribution for the central $F$ matrix
A $(p\times p)$ random symmetric positive definite matrix has a matrix variate beta type II distribution with parameters $(a,b)$ if it has the probability density function
$\frac{\Gamma_{p}(a+b)}{\Gamma_{p}(a)\Gamma_{p}(b)}{\rm det}({\bf V})^{a-(p+1)/2}{\rm det}({\bf I}_{p}+{\bf V})^{-(a+b)},{\bf V}>0,$ (70)
where a>(p1)/2a>(p-1)/2 and b>(p1)/2b>(p-1)/2. We write ${\bf V}\sim B_{p}^{II}(a,b)$ to denote this distribution, which is also known as the matrix variate $F$ distribution. Let $\mathcal{B}_{p}(a,b)$ denote the multivariate beta function
$\mathcal{B}_{p}(a,b)=\frac{\Gamma_{p}(a)\Gamma_{p}(b)}{\Gamma_{p}(a+b)}.$ (71)
Consider the random variable ${\bf V}\sim B_{p}^{II}(n_{1}/2,n_{2}/2)$ and the transformation ${\bf V}={\bf H}\bm{\Lambda}{\bf H}^{\prime}$ from $\bf V$} to its eigenvalues $\bm{\Lambda}={\rm diag}(\lambda_{1},\ldots,\lambda_{p})$ and eigenvectors $\bf H$}, where ${\bf H}\in O(p)$ is in the orthogonal group with the $j$-th column being the normalized eigenvector of $\bf V$} corresponding to the eigenvalue $\lambda_{j}$. The joint distribution of the $p$ eigenvalues $\bm{\Lambda}$ of $\bf V$} is (see [58], Theorem 3.2.17, pp. 104)
$\pi_{\Lambda}(\lambda_{1},\ldots,\lambda_{p})$ $=$ $\frac{\pi^{p^{2}/2}}{\Gamma_{p}(p/2)}\prod_{i<j}|\lambda_{i}-\lambda_{j}|\int_{O(p)}f({\bf H}\bm{\Lambda}{\bf H}^{\prime})(d{\bf H}).$ (72)
The integral can be evaluated as follows
$\int_{O(p)}f({\bf H}\bm{\lambda}{\bf H}^{\prime})(d{\bf H})$ $=$ $\frac{1}{\mathcal{B}_{p}(n_{1}/2,n_{2}/2)}\prod_{j=1}^{p}\lambda_{j}^{(n_{1}-p-1)/2}(1+\lambda_{j})^{-(n_{1}+n_{2})/2}\int_{O(p)}(d{\bf H})$ (73)
$=$ $\frac{1}{\mathcal{B}_{p}(n_{1}/2,n_{2}/2)}\prod_{j=1}^{p}\lambda_{j}^{(n_{1}-p-1)/2}(1+\lambda_{j})^{-(n_{1}+n_{2})/2}$
where (see [58], pp. 104)
$\int_{O(p)}(d{\bf H})=1.$ (74)

### v4

@@
   • We show that the MML threshold for detecting a latent factor agrees with the Baik-Ben Arous-Péché (BBP) phase transition threshold [25].
-The MML threshold is slightly higher than the theoretical distinguishability limit to prevent false positives caused by finite-sample fluctuations of the residual variance estimate.
+MATLAB code implementing the methodology discussed in this paper that can be used to reproduce all experiments in Section IV is available in our github repository.
 II Maximum likelihood estimation
@@
 and $r(\cdot)$ is the marginal distribution of the data
-$r({\bf x})=\sum_{\theta\in\Theta}\pi(\theta)\,p({\bf x}|\theta).$ (14)
+$r({\bf x})=\int_{\theta\in\Theta}\pi(\theta)\,p({\bf x}|\theta)d\theta.$ (14)
 The volume of a cell, $q(C)$, is the coding probability of stating the estimate $\hat{\bm{\theta}}(C)$ for cell $C\in P$, while the estimate used for cell $C$ is obtained by minimising the expected negative log-likelihood over data ${\bf x}\in C$.
@@
 The complete prior distribution over all model parameters is
-$\pi(\bm{\alpha},\sigma,\bm{\phi})=\pi_{\sigma}(\sigma)\,\pi_{\bm{\alpha}}(\alpha_{1},\ldots,\alpha_{J})|J_{{\bf A}\to\bm{\phi}}|\,J!,$ (41)
+$\pi(\bm{\alpha},\sigma,\bm{\phi})=\pi_{\sigma}(\sigma)\,\pi_{\bm{\alpha}}(\alpha_{1},\ldots,\alpha_{J})\pi_{H}({\bf H}_{B})|J_{{\bf A}\to\bm{\phi}}|\,J!,$ (41)
 where the term $J!$ is included because the labelling of the latent factors is arbitrary and $|J_{{\bf A}\to\bm{\phi}}|$ is the transformation of measure from the matrix parametrization $\bf A$} to the orthogonality-preserving parameterization based on Givens rotations.
+Remark.
+In the classical regime ($N\to\infty$, $K$ fixed), finite‑sample PCA eigenvectors concentrate around the population eigenvectors at a rate governed by $N$ and the spike signal-to-noise ratio $\rho_{j}$ $(j\leq J)$.
+Following the perturbation analysis in [44], the principal angle $\theta_{j}$ between the sample PCA eigenvector and the true (population) eigenvector satisfies
+$\mathbb{E}\left\{\sin^{2}\theta_{j}\right\}\asymp\frac{K-1}{N\rho_{j}},\quad\rho_{j}=\alpha_{j}^{2}/\sigma^{2}.$
+This may be incorporated this into the orientation prior $\pi_{H}(\cdot)$ by replacing the uniform Stiefel prior with a matrix Langevin prior conditional on an unknown mean orientation $H_{0}$, with concentration $\kappa_{j}=N\rho_{j}$:
+$\pi_{H}({\bf H}_{B}\mid\alpha,\tau)=c_{K,J}(\bm{\kappa})\,\exp\!\Big\{\mathrm{tr}\!\big(\bm{\kappa}\,{\bf H}_{0}^{\prime}{\bf H}_{B}\big)\Big\},\qquad{\bf H}_{0}\sim\mathrm{Unif}\!\left(V_{J}(\mathbb{R}^{K})\right),$ (42)
+where $H_{0}$ is the unobserved mean orientation with a uniform prior and $c_{K,J}(\bm{\kappa})$ is a normalising constant.
+This approach preserves rotational invariance of the prior while encoding the finite‑sample resolution of the PCA directions.
 III-D Codelength
@@
 $h(\tau)=2\tau^{2}\left(\frac{d\mathcal{I}(\tau)}{d\tau}\right)=L(\tau)-R(\tau),\quad L(\tau)=A\tau-B,\quad R(\tau)=C\tau^{2}\sum_{j=1}^{J}(\delta_{j}-\tau)^{-1},$ (54)
-where $A=N(K-J)-KJ,B=N(K-J)\hat{\tau}_{\rm ML}$ and C=KJ+1C=K-J+1. The solution space of the MML estimate exhibits a phase-transition behavior that can be classified into two regimes:
+where $A=N(K-J)-KJ,B=N(K-J)\hat{\tau}_{\rm ML}$ and C=KJ+1C=K-J+1. The solution space of the MML estimate of $\tau$ exhibits a phase-transition behavior that can be classified into two regimes:
   1. 1.
@@
 Assuming fixed $K,J$ and $N\to\infty$, the bias of the MML estimate of residual variance $\tau:=\sigma^{2}$ is:
-$\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}=\frac{\tau}{N(K-J)}\left(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2}).$ (55)
+$\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}=\frac{\tau}{N(K-J)}\left(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}\right)+\frac{\tau}{N^{2}(K-J)}\mathcal{B}(K,J,\{\rho_{j}\})+O(N^{-3})$ (56)
+where $\mathcal{B}:=\mathcal{B}(K,J,\{\rho_{j}\})$ is
+$\mathcal{B}=-\Big(J+\sum_{m=1}^{J}\rho_{m}^{-1}\Big)\Big[KJ+2(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-2}\Big]-(K-J+1)\sum_{j=1}^{J}\frac{(b_{j}/\tau)}{\rho_{j}^{2}},$ (57)
+and $b_{j}$ is the $O(N^{-1})$ finite-sample mean-shift coefficient of the $j$-th signal eigenvalue, given explicitly by
+$\frac{b_{j}}{\tau}=(1+\rho_{j})\left[\sum_{\ell\leq J,\ \ell\neq j}\frac{1+\rho_{\ell}}{\rho_{j}-\rho_{\ell}}+\frac{K-J}{\rho_{j}}\right].$ (58)
 Proof.
-The derivation uses first-order perturbation theory around the maximum likelihood root.
-All expectations are taken under fixed $K$, $J$ and $N\to\infty$, and higher-order covariance terms between the sample eigenvalues and $\hat{\tau}_{\rm ML}$ are $O(N^{-2})$ and therefore neglected.
-The MML estimate of $\tau$ is a stationary point of the polynomial
+The derivation uses first-order perturbation theory around the maximum likelihood root. All expectations are taken under fixed $K$, $J$ and $N\to\infty$ From [45] and [44], we have for finite NN
+$\delta_{j}=\lambda_{j}+\frac{b_{j}}{N}+O(N^{-2}),\quad j\leq J,$
+where $\lambda_{j}$ is the corresponding population signal eigenvalue and $b_{j}$ $(j\leq J)$ is the $O(N^{-1})$ bias coefficient defined in (58).
+The expected value of the sample signal eigenvalue $\delta_{j}$ [44] is
+$\mathbb{E}\left\{\delta_{j}\right\}$ $=\lambda_{j}+\frac{1}{N}\tau(1+\rho_{j})\left[\sum_{\ell\leq J,\ \ell\neq j}\frac{1+\rho_{\ell}}{\rho_{j}-\rho_{\ell}}+\frac{K-J}{\rho_{j}}\right]=\lambda_{j}+\frac{b_{j}}{N}+O(N^{-2}),\quad j\leq J.$
+Subsequently, the maximum likelihood estimate of $\tau$ can be written as
+$\hat{\tau}_{\rm ML}=\tau+\frac{u}{N}+O(N^{-2}),\quad u=-\tau\sum_{j=1}^{J}(1+\rho_{j}^{-1}),$ (59)
+and the the finite-sample bias of the MLE of $\tau$ (see also [46]) is
+$\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}=\frac{u}{N}+O(N^{-2})=-\frac{\tau}{N}\sum_{j=1}^{J}\left(1+\rho_{j}^{-1}\right)+O(N^{-2}),$ (60)
+up to second order. Recall that the MML estimate of $\tau$ is a stationary point of the polynomial
 $P(\tau)=\sum_{j=1}^{J}a_{j}\tau^{j}=0,\qquad a_{j}=(-1)^{j+1}\left(\hat{\tau}_{\rm ML}e_{J-j}+c_{j}e_{J-j+1}\right),$
@@
 For finite $N$, we will approximate the MML estimate of $\tau$ as the MLE estimate plus a small correction term $\Delta$.
-The bias of the maximum likelihood estimate of $\tau$ up to second order [44] is
-$\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}=-\frac{\tau}{N}\sum_{j=1}^{J}\left(1+\rho_{j}^{-1}\right)+O(N^{-2}).$ (56)
 Let $A(\tau)$ denote the characteristic polynomial of the $J$ sample eigenvalues
@@
 Let k=j1k=j-1 and write the polynomial $Q(\tau)$ as
-$Q(\tau)$ $=\sum_{k=0}^{J}(-1)^{k+2}\left(\frac{k(K-J+1)-KJ}{N(K-J)}\right)e_{J-k}\tau^{k+1}$
-$=\frac{\tau}{N(K-J)}\sum_{k=0}^{J}(-1)^{k}\left(k(K-J+1)-KJ\right)e_{J-k}\tau^{k}$
-$=\frac{\tau}{N(K-J)}\left((K-J+1)\tau A^{\prime}(\tau)-KJA(\tau)\right).$
+$Q(\tau)$ $=\sum_{k=0}^{J}(-1)^{k+2}\left(\frac{k(K-J+1)-KJ}{N(K-J)}\right)e_{J-k}\tau^{k+1}=\frac{\tau}{N(K-J)}\left((K-J+1)\tau A^{\prime}(\tau)-KJA(\tau)\right).$
 Substituting the above into our equation for $\Delta$, we get
-$\Delta$ $=-\frac{\frac{\tau_{\rm ML}}{N(K-J)}\left((K-J+1)\tau A^{\prime}(\tau_{\rm ML})-KJA(\tau_{\rm ML})\right)}{A(\hat{\tau}_{\rm ML})}$
-$=\frac{\hat{\tau}_{\rm ML}}{N(K-J)}\left(KJ-(K-J+1)\sum_{j=1}^{J}\frac{\hat{\tau}_{\rm ML}}{\hat{\tau}_{\rm ML}-\delta_{j}}\right).$
-Next, replace sample quantities with their population counterparts and compute the expectation
-$\mathbb{E}\{\Delta\}$ $=\frac{\tau}{N(K-J)}\left(KJ+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}\right).$
-Recall that the MML estimate of $\tau$ is modelled as the MLE plus a small correction. We finally combine expectations and simplify to obtain the bias estimate up to second order:
-$\mathbb{E}\{\hat{\tau}_{\rm MML}\}$ $=\mathbb{E}\{\hat{\tau}_{\rm MLE}\}+\mathbb{E}\{\Delta\}$
-$=\tau-\frac{\tau}{N}\sum_{j=1}^{J}\left(1+\rho_{j}^{-1}\right)+\frac{\tau}{N(K-J)}\left(KJ+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2})$
-$=\tau+\frac{\tau}{N(K-J)}\left(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}\right)+O(N^{-2}).$
-The maximum likelihood estimate is negatively biased and underestimates noise, with bias of order $O(1/N)$ and approximately proportional to $J\tau/N$.
-In contrast, the MML estimate of $\tau$ is positively biased and overestimates noise in finite samples, with bias of order $O(1/(NK))$ and approximately proportional to $J^{2}\tau/(NK)$.
+$\Delta$ $=\frac{\hat{\tau}_{\rm ML}}{N(K-J)}\left(KJ-(K-J+1)\sum_{j=1}^{J}\frac{\hat{\tau}_{\rm ML}}{\hat{\tau}_{\rm ML}-\delta_{j}}\right),$
+up to $O(N^{-2})$ remainder terms. A first‑order Taylor expansion in $1/N$ gives
+$T=\sum_{j=1}^{J}\frac{\hat{\tau}_{\rm ML}}{\hat{\tau}_{\rm ML}-\delta_{j}}$ $=-\sum_{j=1}^{J}\rho_{j}^{-1}+\frac{1}{N\tau}\sum_{j=1}^{J}\left[-\frac{u}{\rho_{j}}-\frac{u-b_{j}}{\rho_{j}^{2}}\right]+O(N^{-2}),$
+where $u=-\tau\sum_{m=1}^{J}(1+\rho_{m}^{-1})$, as before. Substituting (59) and $T$ into $\Delta$ and simplifying, we get
+$\mathbb{E}\left\{\Delta\right\}$ $=\frac{\tau}{N(K-J)}\left[KJ+(K-J+1)\sum_{j=1}^{J}\rho_{j}^{-1}\right]+\frac{\tau}{N^{2}(K-J)}\mathcal{B}(K,J,\{\rho_{j}\})+O(N^{-3}),$
+where $\mathcal{B}$ is given in (57). Recall that the MML estimate of $\tau$ is modelled as the MLE plus a small correction
+$\mathbb{E}\{\hat{\tau}_{\rm MML}\}$ $=\mathbb{E}\{\hat{\tau}_{\rm MLE}\}+\mathbb{E}\{\Delta\}.$
+Substituting $\mathbb{E}\left\{\Delta\right\}$ and (60) into the above expression and simplifying completes the proof. ∎
+Up to second order, the maximum likelihood estimate is negatively biased and underestimates noise, with bias of order $O(1/N)$ approximately proportional to $J\tau/N$.
+In contrast, the MML estimate of $\tau$ is positively biased and overestimates noise in finite samples, with bias of order $O(1/(NK))$ approximately proportional to $J^{2}\tau/(NK)$.
 The absolute ratio $R$ of the two biases is:
-$R=\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-J)\left(\frac{J+\sum_{j=1}^{J}\rho_{j}^{-1}}{J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}}\right).$ (60)
+$R_{N}=\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-J)\left(\frac{J+\sum_{j=1}^{J}\rho_{j}^{-1}}{(J^{2}+\sum_{j=1}^{J}\rho_{j}^{-1})+\mathcal{B}/N}\right).$ (64)
 Observe that the MML estimate reduces bias compared to the maximum likelihood estimate by a factor roughly proportional to the dimension $K$.
-If the signals are strong (high signal-to-noise ratio) with $\rho_{j}\to\infty$ ($j=1,\ldots,J)$, the ratio is approximately $R\approx(K-J)/J$, suggesting that the MML estimate reduces bias by a factor proportional to the ratio of the total dimension to the latent dimension.
+If the signals are strong (high signal-to-noise ratio) with $\rho_{j}\to\infty$ ($j=1,\ldots,J)$, the ratio $R_{N}$ is
+$R_{N}=\frac{K-J}{J}\left(1+\frac{K}{N}+O(N^{-2})\right),$ (65)
+which is approximately $R\approx(K-J)/J$.
+This suggests that the MML estimate reduces bias by a factor proportional to the ratio of the total dimension to the latent dimension.
 In contrast, for weak signals with $\rho_{j}\to 0$, the terms $\rho_{j}^{-1}$ dominate and the bias reduction factor for the MML estimate is $R\approx(K-J)$.
-Thus, even in the case of weak signals, where the bias of both MML and maximum likelihood estimate is inflated, the relative advantage of the MML estimate remains constant.
+However, in this regime the $J$-factor MML stationary point typically ceases to exist and the MML procedure rejects the $J$-factor model in favor of J1J-1.
 Next, we analyse model selection properties of the MML codelength.
@@
 The absolute bias ratio of the maximum likelihood estimate of $\tau$ to the MML estimate simplifies to
-$\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-1)\left(\frac{1+\sum_{j=1}^{J}\rho_{j}^{-1}}{1^{2}+\sum_{j=1}^{J}\rho_{j}^{-1}}\right)=K-1.$ (68)
-For one true latent factor, the MML estimate reduces bias by the factor of (K1)(K-1) irrespective of the signal-to-noise ratio.
+$\left|\frac{\mathbb{E}\{\hat{\tau}_{\rm ML}-\tau\}}{\mathbb{E}\{\hat{\tau}_{\rm MML}-\tau\}}\right|=(K-1)\frac{1}{1-\frac{K(K+\rho_{1}(\rho_{1}+2))}{N\rho_{1}^{2}}}+O(N^{-2}),$ (73)
+where $\rho_{1}=\alpha_{1}^{2}/\sigma^{2}$ is the signal-to-noise ratio for the top factor.
+As $\rho_{1}\to\infty$, the MML estimate of $\tau$ reduces bias by (K1)/(1K/N)(K-1)/(1-K/N) (to second order).
 When the signal is very weak, both biases blow up in magnitude because the spike is barely separated from the noise bulk.
@@
 For example, we could use the MML PCA codelength in the leaves of a decision tree, similar to the Max-Cut model in [56] or within finite mixture models of probabilistic principal component analyzers, similar to [57, 58, 59].
+VI Data Availability
+The data in this article are available in the MML-PCA GitHub repository at: https://github.com/EnesMakalic/MML-PCA.
 A. Joint eigenvalue distribution for the central $F$ matrix